OpenStudy (anonymous):

A wooden plate is shaped as a pentagon. The measure of one angle is 80°. The remaining interior angles are obtuse and of equal measure. What is the measure of each obtuse angle in the pentagon? 92° 160° 115° 135°

OpenStudy (anonymous):

pentagon can be divided into 3 triangles. Since each triangle has an angle sum of 180 that makes 540 degrees total. We know that one angle is 80 so that leaves 540-80=460 for the remaining 4 angles. Since 460/4=115, choose that answer.

OpenStudy (campbell_st):

the angles sum of a pentagon is 540 degrees if 1 angle is 80 then the other 4 angles add to 540 - 80 = 460 degrees if they remaining angles are obtuse and equal then 460/4 = 115 degrees